Q1. A random variable is defined as “ the number of heads when
a coin tossed”,
(a)twice

(b)three times

(c)four times

(d)five times

Find the PDF of each of them and their respective mean and variance.
1
Expert's answer
2021-08-10T11:51:49-0400

When fair coin is tossed,

The probability to get head = o.5

Probability to get tail =0.5

number of tossed count n = 2, 3, 4, 5

mean = np "=2\times 0.5 = 1"

n= 3

"3\times 0.5 = 1.5"

n= 4

"4\times 0.5 = 2.0"

n=5

"5\times 0.5 = 2.5"

variance "=n\times p\times (1-p)"

n=2

"=2\times 0.5 \times 0.5 = 0.5"

n=3

"=3\times 0.5 \times 0.5 = 0.75"

n=4

"=4\times 0.5\times 0.5 = 2.0"

n=5

"=5\times 0.5\times 0.5 = 1.25"